A 120 line segment is divided into two parts having the ratio 3:5. What are the measures of each part?
Answer by stanbon(75887)

A 120 line segment is divided into two parts having the ratio 3:5. What are the measures of each part?
-----
Equation:
3x + 5x = 120
8x = 120
x = 15
----
3x = 3*15 = 45
5x = 5*15 = 75
====================
Cheers,
Stan H.
================